OpenStudy (anonymous):

Lindsay sells machines which comes in standard and enhanced the standard model sells for $210 and the enhanced sell for $265 if Lindsay sold a total of 19 units and took in $4,485 how many of the standard model did she sell

OpenStudy (anonymous):

210X + 265(19-X) = 4485 solve for X

OpenStudy (shinalcantara):

let 'x' be the number of standard machines she sold 'y' be the number of enhanced machines she sold knowing that the total number of machines she sold is 19 you'll get the equation as: x + y = 19 -----equation 1 since the total amount is 4485 then, @210 for standard and @265 for enhanced 210x + 265y = 4,485 -----equation 2 from multiply equation 1 with 210 [x+y=19] x210 210x + 210y = 3990 -----equation 1' subtract equation 1' from equation 2 210x + 265y = 4,485 - 210x + 210y = 3990 55y = 2595 y = 9 substitute y=9 to equation 1 x + y = 19 x + 9 =19 x = 19-9 x = 10